If the nominal rate on your car loan is 3.95 percent and the rate of inflation is 5.2 percent then the real rate on the loan is negative.

Select one:

True

False

Harold goes to the grocery store to buy his month's supply of Coke. As he enters the soft drink section, he notices that the price of Pepsi has been reduced by 25 percent. He buys Pepsi instead of Coke. This represents the substitution bias that is a problem in the construction of the CPI.

Select one:

True

False

Sally buys a house in 2014, and finances it with a mortgage that carries an annual interest rate of 5 percent. If inflation in 2014 is 2 percent then Sally is paying a real interest rate of 7 percent.

Select one:

True

False

Marvin is a service manager at a car dealership. In year 2000, he earned a salary of $40,000. Suppose the CPI is 160 in year 2000 and 200 in 2010. If Marvin did not want to experience a reduction in his real income for that ten year period then his nominal income in 2010 must be $50,000 or greater.

Select one:

True

False
Answer #1

Q1. True.

Real rate of interest = Nominal rate of interest – Inflation rate = 3.95-5.2 = -1.25 percent

Q2. True.

CPI considers only fixed goods in a basket which do not consider the introduction of new goods or other alternatives

Q3. False

Real interest rate = 5-2 = 3 percent

Q4. True

Year

Salary

CPI

Real Income = (Salary/CPI)*100

2000

40000

160

25000

2010

50000

200

25000